# Total and Symmetry resolved Entanglement spectra in some Fermionic CFTs from the BCFT approach

### v1

@@ Abstract
In this work, we study the universal total and symmetry-resolved entanglement spectra for a single interval of some $2$d Fermionic CFTs using the Boundary Conformal Field theory (BCFT) approach. In this approach, the partition of Hilbert space is achieved by cutting out discs around the entangling boundary points and imposing boundary conditions preserving the extended symmetry under scrutiny. The reduced density moments are then related to the BCFT partition functions and are also found to be diagonal in the symmetry charge sectors. In particular, we first study the entanglement spectra of massless Dirac fermion and modular invariant Dirac fermion by considering the boundary conditions preserving either the axial or the vector $U(1)$ symmetry. The total entanglement spectra of the modular invariant Dirac fermion are shown to match with the compact boson result at the duality radius, while for the massless Dirac fermion, it is found that the boundary entropy term doesn’t match with the self-dual compact boson. The symmetry-resolved entanglement is found to be the same in all cases, except for the charge spectrum which is dependent on both the symmetry and the theory. We also study the entanglement spectra of $N$ massless Dirac fermions by considering boundary conditions preserving different chiral $U(1)^{N}$ symmetries. Entanglement spectra are studied for $U(1)^{M}$ subgroups, where $M\leq N$, by imposing boundary conditions preserving different chiral symmetries. The total entanglement spectra are found to be sensitive to the representations of the $U(1)^{M}$ symmetry in the boundary theory among other behaviours at $O(1)$. Similar results are also found for the Symmetry resolved entanglement entropies. The characteristic $\log\log\left(\ell/\epsilon\right)$ term of the $U(1)$ symmetry is found to be proportional to $M$ in the symmetry-resolved entanglement spectra.
@@ 1 Introduction
The study of the entanglement phenomenon in quantum theories has proven to be one of the most rewarding programs in recent times. It has provided new insights in many fields of physics, perhaps most notable among these are quantum information theory nielsen2010quantum , gauge-gravity duality ryu2006holographic ; ryu2006aspects ; nishioka2009holographic , Black hole physics solodukhin2011entanglement , and quantum many-body systems amico2008entanglement ; vidal2003entanglement . The entanglement spectrum in low dimensional critical systems is sensitive to the corresponding CFT data calabrese2004entanglement ; calabrese2009entanglement2 ; coser2014renyi .
@@
where $\rho_{A}$ is the reduced density matrix. It is given by the trace of the density matrix $\rho$ of the system over degrees of freedom in the subsystem $B$, mathematically denoted by $\rho_{A}=\mathrm{Tr}_{B}\rho$. We mention here that these measures of entanglement fail when we are dealing with mixed-state density matrices since they fail to distinguish the quantum correlations for the classical ones. In the present work, we will only be dealing with the pure states.
@@
In such entanglement studies, especially when the theory is in the continuum, there arises a subtlety while defining the partition. In quantum field theories, the decomposition of the Hilbert space $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ cannot be defined by a sharp cutting operation ohmori2015physics ; witten2018aps . One solution to this problem is to cut our small discs around the boundaries and impose appropriate boundary conditions on the discs ohmori2015physics ; cardy2016entanglement . This is particularly interesting in conformal field theories, as it leads to the study of new universal contributions to the entanglement spectrum when the imposed boundary conditions are conformally invariant. We will discuss this procedure further, relevant to our present program, in Section 2. Such contributions have been studied in the quantum many-body systems in ohmori2015physics ; cardy2010unusual ; hsu2009universal ; cardy2014universal .
@@
In the present work, we will be interested in total and symmetry-resolved entanglement spectrum of some $2$d Fermion conformal theories in the context of Hilbert space decomposition mentioned above. In di2023boundary , it was noticed that when the boundary conditions imposed on the discs preserve the symmetry under scrutiny, it is possible to study all universal contributions to the symmetry-resolved entanglement entropies. The specific case of compact boson was considered in the Reference above. We will consider massless Dirac fermions and modular invariant Dirac fermions for these studies first. Both the theories possess $U(1)$ symmetries, with respect to which we study the symmetry resolution. We also compare the results with the boson case motivated by the Bose-Fermi duality in 2d CFT. We then consider $N$ massless Dirac fermions and consider the $U(1)^{M}$ chiral symmetries studied in smith2020boundary , where $M\leq N$, in the same context. These symmetries are subgroups of the total symmetry group $SO(2N)\times SO(2N)$, thus allowing to study the symmetry resolution of entanglement for smaller subgroups of the symmetry group.
@@ 2.1 Symmetry resolved Entanglement
We consider a theory with global internal symmetry and take the bipartition of the system into subsystem $A$, and its complement $B$. We assume that the theory is in state $\rho$, such that $\left[\rho,\hat{Q}\right]=0$, where $\hat{Q}$ is the charge corresponding to the global internal symmetry. if the charge is additive in subsystems $A$, and $B$, then we have $\left[\rho_{A},\hat{Q}_{A}\right]=0$. This implies that that $\rho_{A}$ is block diagonal in local charge sectors $q$ in subsystem A. This allows us to study entanglement in the local charge sectors. Such studies are called symmetry-resolved entanglement. For this purpose, we define bonsignori2019symmetry
@@
where $\Pi_{q}$ is a unitary projector onto the local charge sector $q$. Note that the trace of $\rho_{A,q}$ is unity. The quantity $\mathrm{Tr}\left(\Pi_{q}\rho_{A}\right)$ is just the probability of obtaining the charge $q$ upon measurement, let us denote this quantity by $p_{q}$ from here on. Then the symmetry resolved Rényi entropies $S_{n,q}$ are given by
@@ 2.2 BCFT approach
To study entanglement for a bipartite, one begins with the decomposition of the Hilbert space $\mathcal{H}$ in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ (or at least the decomposition of the Hilbert space sector in which the state of the system lies). This is not a straightforward exercise in field theories for spatial entanglement, since the Hilbert space $\mathcal{H}$ is built by acting the smeared fields on the vacuum state $\left|\Omega\right.\rangle$. This is done to to ensure that the states thus created have finite norms. So a simple sharp cutting of space into spatial regions $A$, and $B$ doesn’t give a locally well-defined decomposition of the Hilbert space $\mathcal{H}$ ohmori2015physics ; witten2018aps . As discussed in cardy2016entanglement , a well-defined procedure is to consider the projection of the Hilbert space $\mathcal{H}$ onto the space spanned by a complete set of commuting local operators in some open set containing the boundary between the spatial regions $A$, and $B$. In path integral language this is equivalent to imposing the boundary conditions on the boundary of the open set and cutting out the open set from the geometry. This cut in space is characterised by a length parameter $\epsilon$. The parameter $\epsilon$ acts to regularise the entanglement measures.
@@
$Z_{n}^{\alpha\beta}=\mathrm{Tr}_{\mathcal{H}_{\alpha\beta}}\left[q^{\frac{n}{2}\left(L_{0}+\bar{L}_{0}-\frac{c}{12}\right)}\right],$ (8)
where $q=e^{i2\pi\tau}$, with the modular parameter $\tau=i\frac{2\pi}{W}$, and the width of the annulus $W=2\log\frac{\ell}{\epsilon}$. The operators $L_{n}$ are the Virasoro generators of the CFT. The subscript $\mathcal{H}_{\alpha\beta}$ in eq.(8) is to denote that the trace is over the Hilbert space $\mathcal{H}_{\alpha\beta}$. The Hilbert space $\mathcal{H}_{\alpha\beta}$ is usually a priori unknown, however, we may exploit the open/closed string duality to evaluate eq.(8). Using this duality, that is the S-transform of eq.(8), the partition function $Z_{n}$ equals the matrix element
@@
$Z_{n}^{\alpha\beta}=\langle\left.\alpha\right|\tilde{q}^{\frac{n}{2}\left(L_{0}+\bar{L}_{0}-\frac{c}{12}\right)}\left|\beta\right.\rangle,$ (9)
@@
where $\chi_{j}(q)$ is just the trace over the Hilbert space sector $\mathcal{H}_{j}$ and the non-negative integers $n_{j}$ counts the multiplicity of $\mathcal{H}_{j}$ in $\mathcal{H}_{\alpha\beta}$. So, $Z_{n}$ is already diagonalised in the local charge sectors $j$. Now if the conformal boundary conditions $\alpha$ and $\beta$ only preserve a subgroup of the symmetries preserved by either of the boundary conditions, the Hilbert space $\mathcal{H}_{\alpha\beta}$ still decomposes into the charge sectors corresponding to the preserved symmetries.
@@ 3.1 Dirac Fermion
where the phase $\theta\in\left[0,2\pi\right)$. The notation $V$, and $A$ above stand for Vector, and Axial respectively. The boundary states in the $\mathrm{NS}$ sector are known, see for example smith2020boundary . We have however presented a brief calculation of the boundary states in Appendix B. A few good reviews on the construction of boundary states are recknagel2013boundary ; gaberdiel2001conformal . These boundary states are given by
@@
Let’s first give the boundary partition function for the same type of boundary condition imposed on both sides in eq.(9). The boundary partition function $Z_{n}^{\Delta\theta,AA/VV}=\left\langle V/A,\theta\right|\tilde{q}^{\frac{n}{2}\left(L_{0}+\bar{L}_{0}-\frac{c}{12}\right)}\left|V/A,\theta^{\prime}\right\rangle$, where $\Delta\theta=\theta-\theta^{\prime}$, is the same in both cases, so we drop the superscripts in our notation for what follows. The partition function in then is given by
$Z_{n}^{\left(\Delta\theta\right)}(\tilde{q})=\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{p^{2}/2n}}{\eta\left(-\frac{1}{n\tau}\right)}=\frac{\vartheta_{4}\left(\frac{\Delta\theta}{2\pi},-\frac{1}{n\tau}\right)}{\eta\left(-\frac{1}{n\tau}\right)},$ (26)
@@
$Z_{n}^{\left(\Delta\theta\right)}(q)=\sum_{p\in\mathbb{Z}}\frac{q^{\frac{n}{2}\left(\frac{\Delta\theta}{2\pi}+p\right)^{2}}}{\eta(\tau)}.$ (27)
@@
$\begin{split}S_{1}=&\frac{W}{6}-\sum_{k=1}^{\infty}\left(\log\left(1-\tilde{q}^{k}\right)-\frac{k\log(\tilde{q})\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &\qquad+\left[\log(\tilde{q})\frac{\sum_{p\in\mathbb{Z}}p^{2}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}}{\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}}+\log\left(\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}\right)\right].\end{split}$ (29)
@@
$\begin{split}Z_{1}^{AV}=&\tilde{q}^{-1/24}\prod_{k=0}^{\infty}\left(1+\tilde{q}^{k}\right)^{-1}=\frac{\eta\left(-\frac{1}{\tau}\right)}{\eta\left(-\frac{2}{\tau}\right)}\\ =&\sqrt{2}\frac{\eta\left(\frac{\tau}{2}\right)}{\eta\left(\tau\right)}.\end{split}$ (32)
@@
3.2 Modular invariant Dirac fermion
In this subsection, we consider the modular invariant massless Dirac fermions. These fermions are gauged by the operator $(-1)^{F+\bar{F}}$. In this case, we also need to consider the Ramond sector contribution as well. The calculations in the Neveu-Schwartz sector follow from the last subsection. The Ramond sector calculations are a little more involved due to the presence of the fermion zero modes. We have the following mode expansion of the fermion field $\psi(z)$ in the Ramond sector
@@ 3.2 Modular invariant Dirac fermion
where $B_{1},B_{2}\in\{V,A\}$. To evaluate $(-1)^{F+\bar{F}}$ projection, we must have a representation of $(-1)^{F+\bar{F}}$ in the Ramond sector. Since we have already constructed the action of $J_{0}$, and $\bar{J}_{0}$ on the ground states, we may use the $(-1)^{J_{0}+\bar{J}_{0}}$ representation. The states for both boundary conditions in the Neveu-Schwartz sectors are $(-1)^{F+\bar{F}}$ even states. In the Ramond sector, all the states are $(-1)^{F+\bar{F}}$ even and odd states for the Vector and Axial boundary conditions respectively. We therefore have the following BCFT partition function for the same class of boundary conditions imposed on both boundaries
$\begin{split}Z^{\left(\Delta\theta\right)}_{n,V}(\tilde{q})=&g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{p^{2}}{2n}}}{\eta\left(-\frac{1}{n\tau}\right)}+g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}+\frac{1}{2}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{1}{2n}\left(p^{2}-\frac{1}{4}\right)+\frac{1}{8n}}}{\eta\left(-\frac{1}{n\tau}\right)}\\ =&g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{i\frac{p\Delta\theta}{2}}\frac{\tilde{q}^{\frac{p^{2}}{8n}}}{\eta\left(-\frac{1}{n\tau}\right)}\end{split}$ (46)
$Z^{\left(\Delta\theta\right)}_{n,A}(\tilde{q})=$ $g_{A}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{p^{2}}{2n}}}{\eta\left(-\frac{1}{n\tau}\right)},$ (47)
@@
We would expect the two theories to have the same entanglement spectrum headrick2013bose . The issue of entanglement spectrum between the two theories on the torus has been further discussed in lokhande2015modular ; mukhi2018entanglement . We note form eq.(46)-(47), that the partition function of the two theories matches exactly, see ref.di2023boundary for compact boson results. The Vector boundary condition partition function for the modular invariant Dirac fermion matches with the Neumann boundary condition partition function for the free compact boson and the Axial boundary condition partition function matches with the Dirichlet boundary condition partition function. The boundary operator content of the two theories matches exactly, as one would expect from eq.(19)-(20), and eq.(48). We also note that the Axial boundary condition result matches with the corresponding results of the last subsection. The formal reason for this result is due to the fact that all the states in the Ramond sector are $(-1)^{F+\bar{F}}$ odd, and hence don’t contribute.
@@
$Z^{\left(\Delta\theta\right)}_{n,V}(q)=\sum_{p\in\mathbb{Z}}\frac{q^{\frac{n}{2}\left(2p+\frac{\Delta\theta}{2\pi}\right)^{2}}}{\eta\left(n\tau\right)}.$ (49)
@@
$\begin{split}S_{1}=&\frac{W}{6}+\log\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\left(\log\left(1-\tilde{q}^{k}\right)-\frac{k\log(\tilde{q})\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &\qquad+\left[\log(\tilde{q})\frac{\sum_{p\in\mathbb{Z}}p^{2}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}}{4\sum_{p\in\mathbb{Z}}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}}+\log\left(\sum_{p\in\mathbb{Z}}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}\right)\right].\end{split}$ (51)
@@
$Z_{n}^{\mathrm{AV}}=Z_{n}^{\mathrm{VA}}=\frac{\eta\left(\frac{n\tau}{2}\right)}{\eta(n\tau)},$ (52)
we see that there is no anomaly for $\mathrm{AV}$ or $\mathrm{VA}$ boundary conditions for the modular invariant Dirac fermion as opposed to the results of the last subsection. There also exists a $Z_{2}$ symmetry between the two boundary conditions. The symmetry resolution of entanglement entropy with respect to this symmetry has been studied in di2023boundary for free compact boson, and applies similarly in the present study as well.

### v2

@@ Abstract
In this work, we study the universal total and symmetry-resolved entanglement spectra for a single interval of some $2$d Fermionic CFTs using the Boundary Conformal Field theory (BCFT) approach. In this approach, the partition of Hilbert space is achieved by cutting out discs around the entangling boundary points and imposing boundary conditions preserving the extended symmetry under scrutiny. The reduced density moments are then related to the BCFT partition functions and are also found to be diagonal in the symmetry charge sectors. In particular, we first study the entanglement spectra of massless Dirac fermion and modular invariant $Z_{2}$-gauged Dirac fermion by considering the boundary conditions preserving either the axial or the vector $U(1)$ symmetry. The total entanglement spectra of the modular invariant $Z_{2}$-gauged Dirac fermion are shown to match with the compact boson result at the compactification radius where the Bose-Fermi duality holds, while for the massless Dirac fermion, it is found that the boundary entropy term doesn’t match with the self-dual compact boson. The symmetry-resolved entanglement is found to be the same in all cases, except for the charge spectrum which is dependent on both the symmetry and the theory. We also study the entanglement spectra of $N$ massless Dirac fermions by considering boundary conditions preserving different chiral $U(1)^{N}$ symmetries. Entanglement spectra are studied for $U(1)^{M}$ subgroups, where $M\leq N$, by imposing boundary conditions preserving different chiral symmetries. The total entanglement spectra are found to be sensitive to the representations of the $U(1)^{M}$ symmetry in the boundary theory among other behaviours at $O(1)$. Similar results are also found for the Symmetry resolved entanglement entropies. The characteristic $\log\log\left(\ell/\epsilon\right)$ term of the $U(1)$ symmetry is found to be proportional to $M$ in the symmetry-resolved entanglement spectra.
@@ 1 Introduction
The study of the entanglement phenomenon in quantum theories has proven to be one of the most rewarding programs in recent times. It has provided new insights in many fields of physics, perhaps most notable among these are quantum information theory nielsen2010quantum , gauge-gravity duality ryu2006holographic ; ryu2006aspects ; nishioka2009holographic , Black hole physics solodukhin2011entanglement , and quantum many-body systems amico2008entanglement ; vidal2003entanglement . In particular for low dimensional critical systems, the entanglement spectrum is sensitive to the corresponding CFT data calabrese2004entanglement ; calabrese2009entanglement2 ; coser2014renyi .
@@
where $\rho_{A}$ is the reduced density matrix. It is given by the trace of the density matrix $\rho$ of the system over degrees of freedom in the subsystem $B$, mathematically denoted by $\rho_{A}=\mathrm{Tr}_{B}\rho$. We mention here that these measures of entanglement fail when we are dealing with mixed-state density matrices since they fail to distinguish the quantum correlations from the classical ones. In the present work, we will only be dealing with the pure states.
@@
In such entanglement studies, especially when the theory is in the continuum, there arises a subtlety while defining the partition. In quantum field theories, the decomposition of the Hilbert space $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ cannot be defined by a sharp cutting operation ohmori2015physics ; witten2018aps . One solution to this problem is to cut out small discs around the boundaries and impose appropriate boundary conditions on the discs ohmori2015physics ; cardy2016entanglement . This is particularly interesting in conformal field theories, as it leads to the study of new universal contributions to the entanglement spectrum when the imposed boundary conditions are conformally invariant. We will discuss this procedure further, relevant to our present program, in Section 2. Such contributions have been studied in the quantum many-body systems in ohmori2015physics ; cardy2010unusual ; hsu2009universal ; cardy2014universal .
@@
In the present work, we will be interested in total and symmetry-resolved entanglement spectrum of some $2$d Fermion conformal theories in the context of Hilbert space decomposition mentioned above. In di2023boundary , it was noticed that when the boundary conditions imposed on the discs preserve the symmetry under scrutiny, it is possible to study all universal contributions to the symmetry-resolved entanglement entropies. The specific case of compact boson was considered in the Reference above. We will consider massless Dirac fermions and $Z_{2}$-gauged Dirac fermions for these studies first. Since the $Z_{2}$-gauged Dirac fermions considered are modular invariant, we will refer to them as modular invariant Dirac fermions in this work, see Section 3.2. Both the theories possess $U(1)$ symmetries, with respect to which we study the symmetry resolution. We also compare the results with the boson case motivated by the Bose-Fermi duality in 2d CFT. We then consider $N$ massless Dirac fermions and consider the $U(1)^{M}$ chiral symmetries studied in smith2020boundary , where $M\leq N$, in the same context. These symmetries are subgroups of the total symmetry group $SO(2N)\times SO(2N)$, thus allowing to study the symmetry resolution of entanglement for smaller subgroups of the symmetry group.
@@ 2.1 Symmetry resolved Entanglement
We consider a theory with global internal symmetry and take the bipartition of the system into subsystem $A$, and its complement $B$. We assume that the theory is in a state $\rho$, such that $[\rho,\hat{Q}]=0$, where $\hat{Q}$ is the charge corresponding to the global internal symmetry. If the charge is additive in subsystems $A$, and $B$, then we have $[\rho_{A},\hat{Q}_{A}]=0$. This implies that $\rho_{A}$ is block diagonal in local charge sectors $q$ in subsystem A. This allows us to study entanglement in the local charge sectors. Such studies have been termed symmetry-resolved entanglement. Since we are only interested in pure states in this work, we restrict ourselves to pure density matrices $\rho$ from here onwards. To study entanglement in local charge sectors we define bonsignori2019symmetry
@@
where $\Pi_{q}$ is a unitary projector onto the local charge sector $q$. Note that the trace of $\rho_{A,q}$ is unity. The quantity $\mathrm{Tr}\left(\Pi_{q}\rho_{A}\right)$ is just the probability of obtaining the charge $q$ upon measurement, let us denote this quantity by $p_{q}$ from here on. The symmetry resolved Rényi entropies $S_{n,q}$ are given by
@@ 2.2 BCFT approach
To study entanglement for a bipartition, one begins with the decomposition of the Hilbert space $\mathcal{H}$ in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ (or at least the decomposition of the Hilbert space sector in which the state of the system lies). This is not a straightforward exercise in field theories for spatial entanglement, since the Hilbert space $\mathcal{H}$ is built by acting the smeared fields on the vacuum state $\left|\Omega\right.\rangle$. This is done to ensure that the states thus created have finite norms. So a simple sharp cutting of space into spatial regions $A$, and $B$ doesn’t give a locally well-defined decomposition of the Hilbert space $\mathcal{H}$ ohmori2015physics ; witten2018aps . As discussed in cardy2016entanglement , a well-defined procedure is to consider the projection of states in the Hilbert space $\mathcal{H}$ onto a common eigenstate of a complete set of commuting local operators in some open set containing the boundary between the spatial regions $A$, and $B$. In path integral language this is equivalent to imposing some appropriate boundary conditions on the boundary of the open set and cutting out the open set from the geometry. This cut in space is characterised by a length parameter $\epsilon$. The parameter $\epsilon$ acts to regularise the entanglement measures.
@@
$Z_{n}^{\alpha\beta}=\mathrm{Tr}_{\mathcal{H}_{\alpha\beta}}\left[q^{n\left(L_{0}-\frac{c}{24}\right)}\right],$ (8)
where $q=e^{i2\pi\tau}$, with the modular parameter $\tau=i\frac{\pi}{W}$, and the width of the annulus $W=2\log\frac{\ell}{\epsilon}+O(\epsilon)$. The operators $L_{n}$ are the Virasoro generators of the CFT. The subscript $\mathcal{H}_{\alpha\beta}$ in eq.(8) is to denote that the trace is over the Hilbert space $\mathcal{H}_{\alpha\beta}$. The Hilbert space $\mathcal{H}_{\alpha\beta}$ is usually a priori unknown, however, we may exploit the open/closed string duality to evaluate eq.(8). Using this duality, that is the S-transform of eq.(8), the partition function $Z_{n}$ equals the matrix element
@@
$Z_{n}^{\alpha\beta}=\langle\left.\alpha\right|\tilde{q}^{\frac{1}{2n}\left(L_{0}+\bar{L}_{0}-\frac{c}{12}\right)}\left|\beta\right.\rangle,$ (9)
@@
where $\chi_{j}(q)$ is just the trace over the Hilbert space sector $\mathcal{H}_{j}$ and $n_{j}$ is the non-trivial multiplicity of $\mathcal{H}_{j}$ in $\mathcal{H}_{\alpha\beta}$. This implies that the partition function $Z^{\alpha\beta}_{n}$ is already diagonalised in the local charge sectors $j$. Now if the conformal boundary conditions $\alpha$ and $\beta$ only preserve a subgroup of the symmetries preserved by either of the boundary conditions, the Hilbert space $\mathcal{H}_{\alpha\beta}$ still decomposes into the charge sectors corresponding to the preserved symmetries.
@@ 3.1 Dirac Fermion
where the phase $\theta\in\left[0,2\pi\right)$. The notation $V$, and $A$ above stand for Vector, and Axial respectively. The boundary states in the $\mathrm{NS}$ sector are known, see for example smith2020boundary . We have however presented a brief calculation of the boundary states in Appendix B. A few good reviews on the construction of boundary states are recknagel2013boundary ; gaberdiel2001conformal ; blumenhagen2009introduction . These boundary states are given by
@@
Let’s first give the boundary partition function for the same type of boundary condition imposed on both sides in eq.(9). The boundary partition function $Z_{n}^{\Delta\theta,AA/VV}=\left\langle V/A,\theta\right|\tilde{q}^{\frac{1}{2n}\left(L_{0}+\bar{L}_{0}-\frac{c}{12}\right)}\left|V/A,\theta^{\prime}\right\rangle$, where $\Delta\theta=\theta-\theta^{\prime}$, is the same in both cases, so we drop the superscripts in our notation for what follows. The partition function is then given by
$Z_{n}^{\left(\Delta\theta\right)}=\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{p^{2}/2n}}{\eta\left(-\frac{1}{n\tau}\right)}=\frac{\vartheta_{4}\left(\frac{\Delta\theta}{2\pi},-\frac{1}{n\tau}\right)}{\eta\left(-\frac{1}{n\tau}\right)},$ (26)
@@
$Z_{n}^{\left(\Delta\theta\right)}=\sum_{p\in\mathbb{Z}}\frac{q^{\frac{n}{2}\left(\frac{\Delta\theta}{2\pi}+p\right)^{2}}}{\eta(n\tau)}.$ (27)
@@
$\begin{split}S_{1}=&\frac{W}{6}-\sum_{k=1}^{\infty}\left(\log\left(1-\tilde{q}^{k}\right)-\frac{k\log(\tilde{q})\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &\qquad+\left[\frac{1}{2}\log(\tilde{q})\frac{\sum_{p\in\mathbb{Z}}p^{2}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}}{\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}}+\log\left(\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\tilde{q}^{p^{2}/2}\right)\right].\end{split}$ (29)
@@
$\begin{split}Z_{1}^{AV}=&\tilde{q}^{-1/24}\prod_{k=0}^{\infty}\left(1+\tilde{q}^{k}\right)^{-1}=\frac{\eta\left(-\frac{1}{\tau}\right)}{\eta\left(-\frac{2}{\tau}\right)}\\ =&\sqrt{2}\frac{\eta\left(\tau\right)}{\eta\left(\frac{\tau}{2}\right)}.\end{split}$ (32)
@@
3.2 Modular invariant $Z_{2}$-gauged Dirac fermion
In this subsection, we consider the modular invariant massless Dirac fermions. These fermions are gauged by the operator $(-1)^{F+\bar{F}}$. In this case, we also need to consider the Ramond sector contribution as well. The calculations in the Neveu-Schwarz sector follow from the last subsection. The Ramond sector calculations are a little more involved due to the presence of the fermion zero modes. We have the following mode expansion of the fermion field $\psi(z)$ in the Ramond sector
@@ 3.2 Modular invariant $Z_{2}$-gauged Dirac fermion
where $B_{1},B_{2}\in\{V,A\}$. To evaluate $(-1)^{F+\bar{F}}$ projection, we must have a representation of $(-1)^{F+\bar{F}}$ in the Ramond sector. Since we have already constructed the action of $J_{0}$, and $\bar{J}_{0}$ on the ground states, we may use the $(-1)^{J_{0}+\bar{J}_{0}}$ representation. The states for both boundary conditions in the Neveu-Schwarz sectors are $(-1)^{F+\bar{F}}$ even states. In the Ramond sector, all the states are $(-1)^{F+\bar{F}}$ even and odd states for the Vector and Axial boundary conditions respectively. We therefore have the following BCFT partition function for the same class of boundary conditions imposed on both boundaries
$\begin{split}Z^{\left(\Delta\theta\right)}_{n,V}=&g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{p^{2}}{2n}}}{\eta\left(-\frac{1}{n\tau}\right)}+g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}+\frac{1}{2}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{1}{2n}\left(p^{2}-\frac{1}{4}\right)+\frac{1}{8n}}}{\eta\left(-\frac{1}{n\tau}\right)}\\ =&g_{V}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{i\frac{p\Delta\theta}{2}}\frac{\tilde{q}^{\frac{p^{2}}{8n}}}{\eta\left(-\frac{1}{n\tau}\right)}\end{split}$ (46)
$Z^{\left(\Delta\theta\right)}_{n,A}=$ $g_{A}^{2}\tilde{q}^{-\frac{1}{24n}}\sum_{p\in\mathbb{Z}}e^{ip\Delta\theta}\frac{\tilde{q}^{\frac{p^{2}}{2n}}}{\eta\left(-\frac{1}{n\tau}\right)},$ (47)
@@
We would expect the two theories to have the same entanglement spectrum headrick2013bose . The issue of entanglement spectrum between the two theories on the torus has been further discussed in lokhande2015modular ; mukhi2018entanglement . We note from eq.(46)-(47), that the partition function of the two theories matches exactly, see ref.di2023boundary for compact boson results. The Vector boundary condition partition function for the modular invariant Dirac fermion matches with the Neumann boundary condition partition function for the free compact boson and the Axial boundary condition partition function matches with the Dirichlet boundary condition partition function. The boundary operator content of the two theories matches exactly, as one would expect from eq.(19)-(20), and eq.(48). We also note that the Axial boundary condition result matches with the corresponding results of the last subsection. The formal reason for this result is due to the fact that all the states in the Ramond sector are $(-1)^{F+\bar{F}}$ odd, and hence don’t contribute.
@@
$Z^{\left(\Delta\theta\right)}_{n,V}=\sum_{p\in\mathbb{Z}}\frac{q^{\frac{n}{2}\left(2p+\frac{\Delta\theta}{2\pi}\right)^{2}}}{\eta\left(n\tau\right)}.$ (49)
@@
$\begin{split}S_{1}=&\frac{W}{6}+\log\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\left(\log\left(1-\tilde{q}^{k}\right)-\frac{k\log(\tilde{q})\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &\qquad+\left[\frac{1}{8}\log(\tilde{q})\frac{\sum_{p\in\mathbb{Z}}p^{2}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}}{\sum_{p\in\mathbb{Z}}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}}+\log\left(\sum_{p\in\mathbb{Z}}e^{i\frac{p}{2}\Delta\theta}\tilde{q}^{p^{2}/8}\right)\right].\end{split}$ (51)
@@
$Z_{n}^{\mathrm{AV}}=Z_{n}^{\mathrm{VA}}=\frac{\eta(n\tau)}{\eta\left(\frac{n\tau}{2}\right)},$ (52)
we see that there is no anomaly for $\mathrm{AV}$ or $\mathrm{VA}$ boundary conditions for the modular invariant Dirac fermion as opposed to the results of the last subsection. In the open channel, the mixed boundary conditions imply that the current $J(z)$ admits mode expansion in terms of half integer modes $J_{n}$, where $n\in\{\mathbb{Z}+1/2\}$ blumenhagen2009introduction . Consequently the mixed boundary condition partition function is just a single twisted $U(1)$ character built on a primary with conformal dimension $h=\frac{1}{16}$. There is a $Z_{2}$ symmetry corresponding to the Hilbert space sectors containing states constructed by even and odd number of $J_{-n}$ modes, where $n>0$, respectively. The symmetry resolution of entanglement entropy with respect to this symmetry has been studied in di2023boundary for free compact boson, and applies similarly in the present study as well.